Mrs. Lynch has 21 coins in nickles and dimes. Thier total value is $1.65. How many of each coin is there
Answer by bam878s(77)

let x = number of nickels and y = number of dimes.
then x + y = 21
and
.05x + .10y = 1.65
From the first equation, we can solve ofr either x or y and plug into the second
x + y = 21 => x = 21 - y.
.05(21 - y) + .10y = 1.65
1.05 - .05y + .10y = 1.65
1.05 + .05y = 1.65
.05y = .6
y = 12
Now use this value in the first equation and solve for x
x + 12 = 21 => x = 21 -12 = 9
So she has 9 nickels and 12 dimes
check:
9 nickels = .45 and 12 dimes = 1.20
1.20 + .45 = 1.65